# Retraining-Free Merging of Sparse Mixture-of-Experts via Hierarchical Clustering

### v1

@@ B.1 Mixtral 8x7B
We present the activation frequency analysis of all experts in Mixtral 8x7B (Jiang et al., 2024) using our sampling dataset from C4 (Raffel et al., 2020) and eight language benchmarks. The results provide evidence against using frequency as the sole criterion for determining the number of experts in each layer. The analysis reveals variability in activation frequency across different tasks, highlighting the fact that this metric is not a consistent or reliable indicator for expert selection in task-agnostic settings.
@@
Figure 11: The frequency anslysis of TinyLLaMa-4x1.1B-MoE on RTE, Winogrande and HellaSwag.

### v4

@@ E.1 Mixtral 8x7B
We present the activation frequency analysis of all experts in Mixtral 8x7B (Jiang et al., 2024) using our sampling dataset from C4 (Raffel et al., 2020) and eight language benchmarks. The results provide evidence against using frequency as the sole criterion for determining the number of experts in each layer. The analysis reveals variability in activation frequency across different tasks, highlighting the fact that this metric is not a consistent or reliable indicator for expert selection in task-agnostic settings
@@
Figure 13: The frequency anslysis of TinyLLaMa-4x1.1B-MoE on RTE, Winogrand‘e and HellaSwag.